Adam is 8 years older than eve. 17 years ago Adam was twice older than eve. In how many years will Adams age be twice eves age now?
Answer by ankor@dixie-net.com(22740)

let a = Adam's present age
let e = Eve's present age
:
Write an equation for each statement
:
Adam is 8 years older than eve.
a = e + 8
17 years ago Adam was twice older than eve.
a - 17 = 2(e-17)
a - 17 = 2e - 34
a = e - 34 + 17
a = e - 17
replace a with (e+8)
e + 8 = 2e - 17
8 + 17 = 2e - e
25 = e
Eve is 25 yrs old and obviously Adam is 33 yrs
:
In how many years will Adams age be twice eves age now?
let y = no. of yrs for this to be true
a + y = 2e
Replace their ages
33 + y = 2(25)
y = 50 - 33
y = 17 yrs Adam will be twice as old as Eve now